OpenStudy (anonymous):

im having some problems with my math if someone could help id be eternally greatful 1. What are the solutions of 3x2 = 18x? 3.Identify the solutions of 2x2 + 13x + 6 = 0. 4.Solve x2 − 7x + 12 = 0. 7.What are the x-intercept(s) of the graph of y + 6 = x2 − x? 10.What is true about the solutions of a quadratic equation when the radicand in the quadratic formula is negative? 15.Where does the graph of y = 2x2 − 9x − 5 cross the x-axis? thanks for the help everyone that awnsers

OpenStudy (amistre64):

well, assuming zero is already taken care of ... divide by x

OpenStudy (anonymous):

I will help u one at a time so number 1.... \[3x^2=18x\] \[3x^2-18x=0\] \[3x(x-6)=0\] \[x=0,6\]

OpenStudy (anonymous):

alright, I think I understand that one but I got a negative

OpenStudy (anonymous):

mostly cuz I made a typo but thanks

OpenStudy (anonymous):

amistre what do you mean?

OpenStudy (amistre64):

3x2 = 18x it should be obvious that zero is one solution; assuming that we need another one we can then divide each side by x and solve

OpenStudy (amistre64):

3x2 = 18x 3x = 18

OpenStudy (anonymous):

oh,... lol thanks that makes a lot more sense then what I was trying

OpenStudy (anonymous):

hows about the next one?

OpenStudy (amistre64):

can you factor it? or what methods are available to you?

OpenStudy (anonymous):

x = 1 over 2, x = 6 x = −1 over 2, x = −6 x = −3 over 2, x = −2 x = 3 over 2, x = 2 those are the choices

OpenStudy (amistre64):

if you have choices, then the easiest thing to do is to test them out ...

OpenStudy (amistre64):

try the non fractions

OpenStudy (amistre64):

2x2 + 13x + 6 = 0 2x2 + 13x = -6 we are going to need some negative values in order to get this to work out

OpenStudy (amistre64):

does x=-2 work? of not, then its the other one